5x-6y=3
2x=12y=12

Answer by stanbon(75887)

5x-6y=3
2x+12y=12
----------
Solve each for "y":
6y = 5x-3
y = (5/6)x-(1/2)
----------------------
12y=-2x+12
y = (-1/6)x+1
----------------------
Substitute to solve for "x":
(5/6)x-(1/2) = (-1/6)x+1
x = 3/2
---------------
Substitute to solve for "y":
y = (-1/6)(3/2)+1
y = 3/4
====================
Cheers,
Stan H.